# Quantum Key Leasing for PKE and FHE with a Classical Lessor

### 03 Jan 2024 v3

@@
In this work, we consider the problem of secure key leasing, also known as revocable cryptography (Agarwal et. al. Eurocrypt’ 23, Ananth et. al. TCC’ 23), as a strengthened security notion to its predecessor put forward in Ananth et. al. Eurocrypt’ 21. This problem aims to leverage unclonable nature of quantum information to allow a lessor to lease a quantum key with reusability for evaluating a classical functionality. Later, the lessor can request the lessee to provably delete the key and then the lessee will be completely deprived of the capability to evaluate. In this work, we construct a secure key leasing scheme to lease a decryption key of a (classical) public-key, homomorphic encryption scheme from standard lattice assumptions. Our encryption scheme is exactly identical to the (primal) version of Gentry-Sahai-Waters homomorphic encryption scheme with a carefully chosen public key matrix. We achieve strong form of security where:
@@ 1 Introduction
Quantum information, through principles such as the no-cloning theorem, offers excititng possibilities for which there are no classical counterparts. An active area of research are primitives such as copy protection [Aar09] and quantum money [Wie83], which simply can’t be realized classically.
Copy protection, a task that captures a wide variety of unlconability related tasks, is aimed at leveraging the power of unclonable quantum information to prevent piracy of software functionalities. However, constructions of quantum copy protection have turned out to be notoriously hard – in fact, it is shown to be impossible in general [AP21] and only exist relative to structured oracles [${}^{+}$20] if we aim at a generic scheme secure for all copy-protectable functions. Only a handful of functions are known to be copy-protectable without using oracles in the construction, but either they are built from from very strong, less-standard cryptographic assumptions (e.g. post-quantum iO) [CLLZ21, LLQZ22] or the functionalities are very limited (point functions) [CMP20, ${}^{+}$22].
@@
At a later point in time, the lessor may ask the lessee to return or destroy the key for reasons such as lessee’s subscription expiring or the lessor(client) deciding to use another server. The notion of secure key leasing allows the lessee to “delete" the quantum key and create a deletion certificate which can be verified by the lessor. Once the deletion certificate is sent to the lessor and verified, the security guarantee says that the lessee provably loses the associated capability. This must hold even if the lessee is fully malicious and may make arbitrary attempts to copy the quantum state or learn the underlying classical functionality before the deletion. A number of recent papers have studied the notion of secure key leasing for primitives like PKE, FE, FHE and PRF [APV23, ${}^{+}$23], in the context of a quantum lessor and quantum lessee with quantum communication. While the notion of secure software leasing is interesting on its own, remarkly it has also found an application in the seemingly unrelated context of leakage or intrusion detection [ÇGLR23] where an adversary breaks into a machine to steal a secret key and one would like to detect if such an attack happened.
@@
Secure key leasing is a uniquely quantum phonomenon which is enabled by the no-cloning theorem, clearly impossible in the classical setting. Therefore, some quantum capabilities are necessary to enable it. However, a natural question is: to what extent are quantum capabilities necessary? Do both the client and the server need quantum capabilities? Do we require quantum communication in addition to quantum computation? As discussed in [Shm22, CHV23], even in the future where full scale quantum computer come into the household and every user can maintain quantum memory, setting up large-scale end-to-end quantum communication is still expensive and lossy. Moreover, having a client operating only on classical resources and delegating the quantum work to a server is desirable in almost all sorts of circumstances.
@@
A recent beautiful work of Ananth, Poremba, and Vaikuntanathan [APV23] used the dual Regev cryptosystem to build primitives such as public-key encryption (PKE), fully-homomorphic encryption (FHE), and pseudo random functions (PRFs) with key revocation. While their basic constructions require quantum communication as well as a quantum certificate of deletion, they can extend their constructions to achieve classical deletion certificates [APV23] as well. However a key limitation is their work is that the security only holds either based on an unproven conjecture, or only against an adversarial server (leasee) which outputs a valid deletion certificate with probability negligibly close to 1. Proving security against general adversaries based on standard assumptions was left as an open problem. In addition, in both works [${}^{+}$23, APV23], the client and the server must have both quantum capabilities and quantum communication is also necessary in the key generation process. This leads to the following natural open problem:
@@ 1.1 Our Results
(Informal) Assuming the post-quantum subexponental hardness of Learning-with-Errors, there exists a secure key leasing scheme for PKE and FHE with a completely classical lessor/client.
@@
    The key generation is a two-message protocol with one classical message from the lessor to lessee and one back.
@@
A key contribution of our work is to connect our problem to classically verifiable proof of quantumness [${}^{+}$21]. The latter is a well developed research direction with beautiful ideas. This connection allows us to design modular constructions and write proofs of security that are arguably simpler, thanks to a number of technical ideas that we could rely on the literature in classical verification of quantum computation.
@@
Given a quantum distinguisher with an auxiliary quantum state, that distinguishes LWE samples from random with inverse polynomial probability for some arbitrary secret, there exists an extractor running in time polynomial (in the the secret size, the distinguisher’s time and inverse its advantage) that produces the secret with probability negligibly close to 1.
To the best of our knowledge, this is the first search-to-decision reduction with an almost-perfect search when using quantum algorithms with auxiliary quantum states. To prove the above theorem, we rely on a measurement implementation on the distinguisher’s quantum state [Zha20, ${}^{+}$20] and make new observations about this measurement procedure, by aligning its properties with the features required by the classical LWE search-to-decision reduction by Regev [Reg05a].
We believe this technique and the new observations made maybe of independent interest.
@@ 1.2 Comparison with Related Works
[BBK22] showed a general method to lift classical reductions to post-quantum reductions for a class of primitives satisfying certain constraints. Even though the classical search-to-decision reduction for LWE fit into the framework required in the above lifting theorem, our proof roadmap requires a stronger statement than the one shown in [BBK22].To the best of our knowledge, there is no generic or black-box method to amplify the reduction in [BBK22] for our use. Besides, our techniques differ from [BBK22] in that we do not perform the more involved "state repair" procedure used by [BBK22], which was inspired by [CMSZ22]. Therefore, we believe our techniques may be of independent interest and can shed light on new observations in quantum reductions.
@@ 2 Technical Overview
Our security requirement is also fairly intuitive. In the security game, the challenger generates a classical $\mathsf{mpk}$ along with a trapdoor $\mathsf{td}$ and gives $\mathsf{mpk}$ to a quantum polynomial time attacker $\mathcal{A}$. $\mathcal{A}$ generates a quantum decryption key $\rho_{\mathsf{sk}}$ and its associated public key $\mathsf{pk}$. $\mathcal{A}$ publishes its $\mathsf{pk}$. The challenger then asks $\mathcal{A}$ to provide a certificate $\mathsf{cert}$. The game aborts if $\mathsf{cert}$ does not verify. However, if the certificate verifies, we require that $\mathcal{A}$ should not be able to distinguish an encryption of 00 from an encryption of $1$ with a non-negligible probability. We formalize this by requiring that after the step of successful verification of certificate, an adversary outputs a quantum state $\rho_{\mathsf{Delete}}$ as a adversarial quantum decyrptor.
@@
The definitions are inspired from the framework introduced by Zhandry [Zha20](see Section 8.2,Section 5 for details) and will imply the past PKE-SKL definition used in [${}^{+}$23, APV23]. As we will discuss later, defining the security via Threshold Implementation also assists us in our security proof.
@@
Consider the following game: The challenger samples $\mathbf{A}$ and an LWE sample $\mathbf{k}=\mathbf{s}\mathbf{A}+\mathbf{e}\mod q$ where $\mathbf{s}$ is (say) binary random secret $\mathbf{s}$ and $\mathbf{e}$ is chosen from a discrete Gaussian distribution with parameter $\sigma$. For any such sample, there exists a BQP algorithm that could sample a vector $\mathbf{y}=\mathbf{x}_{0}\mathbf{A}+\mathbf{e}^{\prime}$ where $\mathbf{e}^{\prime}$ chosen from discrete Gaussian with parameter $\sigma^{\prime}\gg\sigma$ (by a superpolynomial factor), along with a special state ${\sf\rho_{\mathsf{sk}}}$. The state is of the form ${\sf\rho_{\mathsf{sk}}}=\ket{(0,\mathbf{x}_{0})}+\ket{(1,\mathbf{x}_{1})}$ where $\mathbf{x}_{1}=\mathbf{s}-\mathbf{x}_{0}$. Measuring this state yields either $\mathbf{x}_{0}$ so that $\mathbf{x}_{0}\mathbf{A}\approx\mathbf{y}$ or $\mathbf{x}_{1}$ so that $\mathbf{x}_{1}\mathbf{A}$ is close to $\mathbf{y}-\mathbf{k}$. On the other hand if we measured in the Hadamard basis we will again obtain $(\mathbf{d},\alpha)$ for a random string $\mathbf{d}$ so that $\alpha=\langle\mathbf{d},\mathbf{x}_{0}\oplus\mathbf{x}_{1}\rangle$  [^2] In the actual protocol, $\alpha$ is not equal to $\langle\mathbf{d},\mathbf{s}\rangle$ where $\mathbf{s}=\mathbf{x}_{0}\oplus\mathbf{x}_{1}$. In fact, we first apply a binary decomposition to $\mathbf{x}_{0},\mathbf{x}_{1}$ and then measure in Hadamard basis. i.e. we have $\alpha=\langle\mathbf{d},\mathcal{J}(\mathbf{x}_{0})\oplus\mathcal{J}(\mathbf{x}_{1})\rangle\mod 2$, where $\mathcal{J}$ is the invertible binary decomposition function. Such a binary decomposition is needed for the security proof of the NTCF to go through in [${}^{+}$21].
@@
Now we gradually approach a scheme similar to the two-key Regev PKE scheme which can benefit from the NTCF security properties described above. The idea is that we have a public key $\mathbf{A}$, an LWE sample $\mathbf{k}=\mathbf{s}\mathbf{A}+\mathbf{e}$ where $\mathbf{s}$ is a random binary secret and $\mathbf{e}$ is Gaussian distributed exactly like the distribution used for proof of quantumness. Along with $\mathbf{k}$, the public key also consists of an LWE sample $\mathbf{y}=\mathbf{x}_{0}\mathbf{A}+\mathbf{e}^{\prime}$ chosen as per the distribution in the proof of quantumness again. The client manintains a trapdoor $\mathsf{td}$: $\mathsf{td}$ consists of a trapdoor matrix $\mathbf{T}$ of $\mathbf{A}$ that can be generated at the same time as sampling $\mathbf{A}$ (recall that $\mathsf{td}$ will be used to verify the deletion certificate later). The leased decryption state $\rho_{\mathsf{sk}}$ is the superposition |$\ket{0,x_{0}}+\ket{1,x_{1}}$ where $\mathbf{x}_{1}=\mathbf{x}_{0}-\mathbf{s}$.
To encrypt a bit $\mu$, one samples a random binary string $\mathbf{r}$ and computes a ciphertext $\mathsf{ct}=(\mathsf{ct}_{1},\mathsf{ct}_{2},\mathsf{ct}_{3})$ where $\mathsf{ct}_{1}=\mathbf{A}\mathbf{r}$, $\mathsf{ct}_{2}=\mathbf{k}\mathbf{r}$ and $\mathsf{ct}_{3}=\mathbf{y}\mathbf{r}+\mu\lceil\frac{q}{2}\rceil$. Observe that the ciphertext can be decrypted by both $\mathbf{x}_{0}$ or $\mathbf{x}_{1}$. This is because $\mathsf{ct}_{3}-\mathbf{x}_{0}\mathbf{A}$ is close to $\mu\lceil\frac{q}{2}\rceil$ and similarly, $-\mathsf{ct}_{2}+\mathsf{ct}_{3}+\mathbf{x}_{1}\mathbf{A}$ is also close to $\mu\lceil\frac{q}{2}\rceil$. Thus, there is a way to decrypt a ciphertext cohenrently without disturbing the state $\rho_{\mathsf{sk}}$, by the gentle measurement lemma [Aar04]. Accordingly, we set the deletion certificate to be the string $(\alpha,\mathbf{d})$ to use the property of NTCF.
@@
We first discuss a very weak definition of security which the above simple scheme (almost) satisfies already. In this definition, a successful BQP adversary provides a deletion certificate that’s valid with probability $1-\mathsf{negl}$ for some negligible $\mathsf{negl}$ and then conditioned on that distinguishes ciphertexts with a noticable probability. One observation is that given a decryptor that’s capable of distinguishing encryptions of zero from encryptions of one, must also enable distinguishing ciphertexts of zero of the form $(\mathsf{ct}_{1},\mathsf{ct}_{2},\mathsf{ct}_{3})$ where $\mathsf{ct}_{1}=\mathbf{A}\cdot\mathbf{r}$, $\mathsf{ct}_{2}=\mathbf{k}\cdot\mathbf{r}$ and $\mathsf{ct}_{3}=\mathbf{y}\cdot\mathbf{r}$ from truly random strings with a noticable probabilty. This means that this distinguisher distinguishes samples of the form $(\mathbf{A}\mathbf{r},\mathbf{k}\mathbf{r},\mathbf{y}\mathbf{r})$ from random. Since $\mathbf{k}$ is pseudorandom due to LWE it must also distinguish $(\mathbf{A}\mathbf{r},\mathbf{k}\mathbf{r},\mathbf{y}\mathbf{r})$ from random where $\mathbf{k}$ is chosen at random. Observe that $\mathbf{y}=\mathbf{x}_{0}\mathbf{A}+\mathbf{e}^{\prime}$, $\mathbf{y}\mathbf{r}=\mathbf{x}_{0}\mathbf{A}\mathbf{r}+\mathbf{e}^{\prime}\mathbf{r}$. Given that $\mathbf{e}^{\prime}\mathbf{r}$ losses information on $\mathbf{r}$, and $\mathbf{A},\mathbf{k}$ are now chosen at random one can appeal to LHL to show that such a distinguisher should distinguish between $(\mathsf{ct}_{1},\mathsf{ct}_{2},\mathsf{ct}_{3})$ generated as $(\mathbf{a},u,\langle\mathbf{x}_{0},\mathbf{a}\rangle+\mathbf{e}^{\prime}\mathbf{r})$ from random, where $\mathbf{a}\in\mathbb{Z}^{n\times 1}_{q}$ and $u\in\mathbb{Z}_{q}$ are random.
Together $(\mathsf{ct}_{1},\mathsf{ct}_{3})$ are now almost distributed as an LWE sample in $\mathbf{x}_{0}$. If we have a quantum search to decision reduction for LWE using the quantum distinguisher (with an internal state), we should be able to extract $\mathbf{x}_{0}$ efficiently giving rise to a contradiction. That is, this reduction would have produced a valid $(\alpha,\mathbf{d})$ with overwhelming probability and conditioned on that $\mathbf{x}_{0}$ with inverse polynomial probability (something ruled out by the NTCF in [${}^{+}$21]).
@@
There is a minor flaw in the above argument that can be fixed. The distribution $(\mathsf{ct}_{1},\mathsf{ct}_{2},\mathsf{ct}_{3})$ formed as $(\mathbf{A}\mathbf{r},\mathbf{k}\mathbf{r},\mathbf{y}\mathbf{r})$ does not behave statistically close to $(\mathbf{a},u,\langle\mathbf{a},\mathbf{x}_{0}\rangle+e)$ for truly random $\mathbf{a},u$ and an error $e$ sampled according to LWE distribution. The reason for that is that if we consider $\mathbf{y}\mathbf{r}=\mathbf{x}_{0}\mathbf{A}\mathbf{r}+\mathbf{e}^{\prime}\mathbf{r}$, the error $e=\mathbf{e}^{\prime}\mathbf{r}$ might not behave as an statistically independent discrete Gaussian. To fix this issue, we modify our encryption algorithm to have a smudging noise $\mathbf{e}^{\prime\prime}$ with superpolynomially larger parameter $\sigma^{\prime\prime}\gg\sigma$ and construct $(\mathsf{ct}_{1},\mathsf{ct}_{2},\mathsf{ct}_{3})$ as $(\mathbf{A}\mathbf{r},\mathbf{k}\mathbf{r},\mathbf{y}\mathbf{r}+\mathbf{e}^{\prime\prime}+\mu\lceil\frac{q}{2}\rceil)$. With this smudging $\mathbf{e}^{\prime}\mathbf{r}$ can now be drowned by $\mathbf{e}^{\prime\prime}$ effectively now behaving as a fresh LWE sample in $\mathbf{x}_{0}$ with slightly larger noise. Accordingly, we design our search-to-decision reduction to also work for samples statistically-close to LWE versus samples close to random, as opposed to the simple case of real LWE versus truly random.

### 06 Mar 2024 v4

@@
In this work, we consider the problem of secure key leasing, also known as revocable cryptography (Agarwal et. al. Eurocrypt’ 23, Ananth et. al. TCC’ 23), as a strengthened security notion to its predecessor put forward in Ananth et. al. (Eurocrypt’ 21). This problem aims to leverage unclonable nature of quantum information to allow a lessor to lease a quantum key with reusability for evaluating a classical functionality. Later, the lessor can request the lessee to provably delete the key and then the lessee will be completely deprived of the capability to evaluate. In this work, we construct a secure key leasing scheme to lease a decryption key of a (classical) public-key, homomorphic encryption scheme from standard lattice assumptions. Our encryption scheme is exactly identical to the (primal) version of Gentry-Sahai-Waters homomorphic encryption scheme with a carefully chosen public key matrix. We achieve strong form of security where:
@@ 1 Introduction
Quantum information, through principles such as the no-cloning theorem, offers exciting possibilities for which there are no classical counterparts. An active area of research are primitives such as copy protection [Aar09] and quantum money [Wie83], which simply can’t be realized classically.
Copy protection, a task that captures a wide variety of unclonability related tasks, is aimed at leveraging the power of unclonable quantum information to prevent piracy of software functionalities. However, constructions of quantum copy protection have turned out to be notoriously hard – in fact, it is shown to be impossible in general [AP21] and only exist relative to structured oracles [${}^{+}$20] if we aim at a generic scheme secure for all copy-protectable functions. Only a handful of functions are known to be copy-protectable without using oracles in the construction, but either they are built from very strong, less-standard cryptographic assumptions (e.g. post-quantum iO) [CLLZ21, LLQZ22] or the functionalities are very limited (point functions) [CMP20, ${}^{+}$22].
@@
At a later point in time, the lessor may ask the lessee to return or destroy the key for reasons such as lessee’s subscription expiring or the lessor(client) deciding to use another server. The notion of secure key leasing allows the lessee to “delete" the quantum key and create a deletion certificate which can be verified by the lessor. Once the deletion certificate is sent to the lessor and verified, the security guarantee says that the lessee provably loses the associated capability. This must hold even if the lessee is fully malicious and may make arbitrary attempts to copy the quantum state or learn the underlying classical functionality before the deletion. A number of recent papers have studied the notion of secure key leasing for primitives like PKE, FE, FHE and PRF [APV23, ${}^{+}$23], in the context of a quantum lessor and quantum lessee with quantum communication. While the notion of secure software leasing is interesting on its own, remarkably it has also found an application in the seemingly unrelated context of leakage or intrusion detection [ÇGLR23] where an adversary breaks into a machine to steal a secret key and one would like to detect if such an attack happened.
@@
Secure key leasing is a uniquely quantum phenomenon which is enabled by the no-cloning theorem, clearly impossible in the classical setting. Therefore, some quantum capabilities are necessary to enable it. However, a natural question is: to what extent are quantum capabilities necessary? Do both the client and the server need quantum capabilities? Do we require quantum communication in addition to quantum computation? As discussed in [Shm22, CHV23], even in the future when full scale quantum computer come into the household and every user can maintain quantum memory, setting up large-scale end-to-end quantum communication is still expensive and lossy. Moreover, having a client operating only on classical resources and delegating the quantum work to a server is desirable in almost all sorts of circumstances.
@@
A recent beautiful work of Ananth, Poremba, and Vaikuntanathan [APV23] used the dual Regev cryptosystem to build primitives such as public-key encryption (PKE), fully-homomorphic encryption (FHE), and pseudo random functions (PRFs) with key revocation. While their basic constructions require quantum communication as well as a quantum certificate of deletion, they can extend their constructions to achieve classical deletion certificates [APV23] as well. However, a key limitation is their work is that the security only holds either based on an unproven conjecture, or only against an adversarial server (leasee) which outputs a valid deletion certificate with probability negligibly close to 1. Proving security against general adversaries based on standard assumptions was left as an open problem. In addition, in both works [${}^{+}$23, APV23], the client and the server must have both quantum capabilities and quantum communication is also necessary in the key generation process. This leads to the following natural open problem:
@@ 1.1 Our Results
(Informal) Assuming the post-quantum sub-exponental hardness of Learning-with-Errors, there exists a secure key leasing scheme for PKE and FHE with a completely classical lessor/client.
@@
    The key generation is a one-message protocol with one classical message from the lessor to lessee and one back.  [^2] As discussed in Appendix A, we need to add a second message to deal with impostor security, which is orthogonal to the major anti-piracy security we prove. For the major anti-piracy security defined in Definition 8.2,Definition 8.6, the protocol suffices to be a one-message protocol from lessor to lessee..
@@
A key contribution of our work is to connect our problem to classically verifiable proof of quantumness [${}^{+}$21]. The latter is a well-developed research direction with beautiful ideas. This connection allows us to design modular constructions and write proofs of security that are arguably simpler, thanks to a number of technical ideas that we could rely on the literature in classical verification of quantum computation.
@@
Given a quantum distinguisher with an auxiliary quantum state, that distinguishes LWE samples from random with inverse polynomial probability for some arbitrary secret, there exists an extractor running in time polynomial (in the secret size, the distinguisher’s time and inverse its advantage) that produces the secret with probability negligibly close to 1.
To the best of our knowledge, this is the first search-to-decision reduction with an almost-perfect search when using quantum algorithms with auxiliary quantum states. To prove the above theorem, we rely on a measurement implementation on the distinguisher’s quantum state [MW05, Zha20, ${}^{+}$20]. In particular, we make new observations about this measurement procedure, by aligning its properties with the features required by the classical LWE search-to-decision reduction by Regev [Reg05a].
We take a markedly distinct approach compared to the previous work [BBK22] on post-quantum reductions, even though both of our approaches set basis upon the [MW05, Zha20] measurement techniques. We believe this technique and the new observations made maybe of independent interest, such as lifting other [Reg05a]-like reductions from classical to post-quantum.
@@ 1.2 Comparison with Related Works
[BBK22] showed a nice generic method to lift classical reductions to post-quantum reductions for a class of primitives satisfying certain constraints. Even though the classical search-to-decision reduction for LWE fit into the framework required in the above lifting theorem, our proof roadmap requires a stronger statement than the one shown in [BBK22].To the best of our knowledge, there is no generic or black-box method to modify the statement in [BBK22] for our use. Besides, our techniques differ significantly from [BBK22] (and [CMSZ22], by which [BBK22] was inspired) in that we do not perform the "nested alternating projections" procedure used in [BBK22, CMSZ22], even though both of our approaches are based on applying the measurement procedures proposed in [MW05, Zha20]. The way we use [MW05, Zha20] aligns intuitively closer with the classical reduction in [Reg05a] and has a simpler proof overall. Therefore, we believe our techniques may be of independent interest and can shed light on new observations in (post-)quantum reductions.
@@ 2 Technical Overview
Our security requirement is also fairly intuitive. In the security game, the challenger generates a classical $\mathsf{mpk}$ along with a trapdoor $\mathsf{td}$ and gives $\mathsf{mpk}$ to a quantum polynomial time attacker $\mathcal{A}$. $\mathcal{A}$ generates a quantum decryption key $\rho_{\mathsf{sk}}$ and its associated public key $\mathsf{pk}$. $\mathcal{A}$ publishes its $\mathsf{pk}$. The challenger then asks $\mathcal{A}$ to provide a certificate $\mathsf{cert}$. The game aborts if $\mathsf{cert}$ does not verify. However, if the certificate verifies, we require that $\mathcal{A}$ should not be able to distinguish an encryption of 00 from an encryption of $1$ with a non-negligible probability. We formalize this by requiring that after the step of successful verification of the certificate, an adversary outputs a quantum state $\rho_{\mathsf{Delete}}$ as a adversarial quantum decyrptor.
@@
The definitions are inspired from the framework introduced by Zhandry [Zha20](see Section 8.2,Section 5 for details) and will imply the existing PKE-SKL definition used in [${}^{+}$23, APV23]. As we will discuss later, defining the security via Threshold Implementation also assists us in our security proof.
@@
Consider the following game: The challenger samples $\mathbf{A}$ and an LWE sample $\mathbf{k}=\mathbf{s}\mathbf{A}+\mathbf{e}\mod q$ where $\mathbf{s}$ is (say) binary random secret $\mathbf{s}$ and $\mathbf{e}$ is chosen from a discrete Gaussian distribution with parameter $\sigma$. For any such sample, there exists a BQP algorithm that could sample a vector $\mathbf{y}=\mathbf{x}_{0}\mathbf{A}+\mathbf{e}^{\prime}$ where $\mathbf{e}^{\prime}$ chosen from discrete Gaussian with parameter $\sigma^{\prime}\gg\sigma$ (by a superpolynomial factor), along with a special state ${\sf\rho_{\mathsf{sk}}}$. The state is of the form ${\sf\rho_{\mathsf{sk}}}=\ket{0,\mathbf{x}_{0}}+\ket{1,\mathbf{x}_{1}}$ where $\mathbf{x}_{1}=\mathbf{s}-\mathbf{x}_{0}$. Measuring this state yields either $\mathbf{x}_{0}$ so that $\mathbf{x}_{0}\mathbf{A}\approx\mathbf{y}$ or $\mathbf{x}_{1}$ so that $\mathbf{x}_{1}\mathbf{A}$ is close to $\mathbf{y}-\mathbf{k}$. On the other hand, if we measured in the Hadamard basis we will again obtain $(\mathbf{d},\alpha)$ for a random string $\mathbf{d}$ so that $\alpha=\langle\mathbf{d},\mathbf{x}_{0}\oplus\mathbf{x}_{1}\rangle$  [^3] In the actual protocol, $\alpha$ is not equal to $\langle\mathbf{d},\mathbf{s}\rangle$ where $\mathbf{s}=\mathbf{x}_{0}\oplus\mathbf{x}_{1}$. In fact, we first apply a binary decomposition to $\mathbf{x}_{0},\mathbf{x}_{1}$ and then measure in Hadamard basis. i.e. we have $\alpha=\langle\mathbf{d},\mathcal{J}(\mathbf{x}_{0})\oplus\mathcal{J}(\mathbf{x}_{1})\rangle\mod 2$, where $\mathcal{J}$ is the invertible binary decomposition function. Such a binary decomposition is needed for the security proof of the NTCF to go through in [${}^{+}$21].
@@
Now we gradually approach a scheme similar to the two-key Regev PKE scheme which can benefit from the NTCF security properties described above. The idea is that we have a public key $\mathbf{A}$, an LWE sample $\mathbf{k}=\mathbf{s}\mathbf{A}+\mathbf{e}$ where $\mathbf{s}$ is a random binary secret and $\mathbf{e}$ is Gaussian distributed exactly like the distribution used for proof of quantumness. Along with $\mathbf{k}$, the public key also consists of an LWE sample $\mathbf{y}=\mathbf{x}_{0}\mathbf{A}+\mathbf{e}^{\prime}$ chosen as per the distribution in the proof of quantumness again. The client maintains a trapdoor $\mathsf{td}$: $\mathsf{td}$ consists of a trapdoor matrix $\mathbf{T}$ of $\mathbf{A}$ that can be generated at the same time as sampling $\mathbf{A}$ (recall that $\mathsf{td}$ will be used to verify the deletion certificate later). The leased decryption state $\rho_{\mathsf{sk}}$ is the superposition |$\ket{0,x_{0}}+\ket{1,x_{1}}$ where $\mathbf{x}_{1}=\mathbf{x}_{0}-\mathbf{s}$.
To encrypt a bit $\mu$, one samples a random binary string $\mathbf{r}$ and computes a ciphertext $\mathsf{ct}=(\mathsf{ct}_{1},\mathsf{ct}_{2},\mathsf{ct}_{3})$ where $\mathsf{ct}_{1}=\mathbf{A}\mathbf{r}$, $\mathsf{ct}_{2}=\mathbf{k}\mathbf{r}$ and $\mathsf{ct}_{3}=\mathbf{y}\mathbf{r}+\mu\lceil\frac{q}{2}\rceil$. Observe that the ciphertext can be decrypted by both $\mathbf{x}_{0}$ or $\mathbf{x}_{1}$. This is because $\mathsf{ct}_{3}-\mathbf{x}_{0}\mathbf{A}$ is close to $\mu\lceil\frac{q}{2}\rceil$ and similarly, $-\mathsf{ct}_{2}+\mathsf{ct}_{3}+\mathbf{x}_{1}\mathbf{A}$ is also close to $\mu\lceil\frac{q}{2}\rceil$. Thus, there is a way to decrypt a ciphertext coherently without disturbing the state $\rho_{\mathsf{sk}}$, by the gentle measurement lemma [Aar04]. Accordingly, we set the deletion certificate to be the string $(\alpha,\mathbf{d})$ to use the property of NTCF.
@@
We first discuss a very weak definition of security which the above simple scheme (almost) satisfies already. In this definition, a successful BQP adversary provides a deletion certificate that’s valid with probability $1-\mathsf{negl}$ for some negligible $\mathsf{negl}$ and then conditioned on that distinguishes ciphertexts with a noticeable probability. One observation is that given a decryptor that’s capable of distinguishing encryptions of zero from encryptions of one, must also enable distinguishing ciphertexts of zero of the form $(\mathsf{ct}_{1},\mathsf{ct}_{2},\mathsf{ct}_{3})$ where $\mathsf{ct}_{1}=\mathbf{A}\cdot\mathbf{r}$, $\mathsf{ct}_{2}=\mathbf{k}\cdot\mathbf{r}$ and $\mathsf{ct}_{3}=\mathbf{y}\cdot\mathbf{r}$ from truly random strings with a noticeable probability. This means that this distinguisher distinguishes samples of the form $(\mathbf{A}\mathbf{r},\mathbf{k}\mathbf{r},\mathbf{y}\mathbf{r})$ from random. Since $\mathbf{k}$ is pseudorandom due to LWE it must also distinguish $(\mathbf{A}\mathbf{r},\mathbf{k}\mathbf{r},\mathbf{y}\mathbf{r})$ from random where $\mathbf{k}$ is chosen at random. Observe that $\mathbf{y}=\mathbf{x}_{0}\mathbf{A}+\mathbf{e}^{\prime}$, $\mathbf{y}\mathbf{r}=\mathbf{x}_{0}\mathbf{A}\mathbf{r}+\mathbf{e}^{\prime}\mathbf{r}$. Given that $\mathbf{e}^{\prime}\mathbf{r}$ losses information on $\mathbf{r}$, and $\mathbf{A},\mathbf{k}$ are now chosen at random, one can appeal to LHL to show that such a distinguisher should distinguish between $(\mathsf{ct}_{1},\mathsf{ct}_{2},\mathsf{ct}_{3})$ generated as $(\mathbf{a},u,\langle\mathbf{x}_{0},\mathbf{a}\rangle+\mathbf{e}^{\prime}\mathbf{r})$ from random, where $\mathbf{a}\in\mathbb{Z}^{n\times 1}_{q}$ and $u\in\mathbb{Z}_{q}$ are random.
Together, $(\mathsf{ct}_{1},\mathsf{ct}_{3})$ are now almost distributed as an LWE sample in $\mathbf{x}_{0}$. If we have a quantum search to decision reduction for LWE using the quantum distinguisher (with an internal state), we should be able to extract $\mathbf{x}_{0}$ efficiently, giving rise to a contradiction. That is, this reduction would have produced a valid $(\alpha,\mathbf{d})$ with overwhelming probability and conditioned on that $\mathbf{x}_{0}$ with inverse polynomial probability (something ruled out by the NTCF in [${}^{+}$21]).
@@
There is a minor flaw in the above argument that can be fixed. The distribution $(\mathsf{ct}_{1},\mathsf{ct}_{2},\mathsf{ct}_{3})$ formed as $(\mathbf{A}\mathbf{r},\mathbf{k}\mathbf{r},\mathbf{y}\mathbf{r})$ does not behave statistically close to $(\mathbf{a},u,\langle\mathbf{a},\mathbf{x}_{0}\rangle+e)$ for truly random $\mathbf{a},u$ and an error $e$ sampled according to LWE distribution. The reason for that is that if we consider $\mathbf{y}\mathbf{r}=\mathbf{x}_{0}\mathbf{A}\mathbf{r}+\mathbf{e}^{\prime}\mathbf{r}$, the error $e=\mathbf{e}^{\prime}\mathbf{r}$ might not behave as an statistically independent discrete Gaussian. To fix this issue, we modify our encryption algorithm to have a smudging noise $\mathbf{e}^{\prime\prime}$ with superpolynomially larger parameter $\sigma^{\prime\prime}\gg\sigma$ and construct $(\mathsf{ct}_{1},\mathsf{ct}_{2},\mathsf{ct}_{3})$ as $(\mathbf{A}\mathbf{r},\mathbf{k}\mathbf{r},\mathbf{y}\mathbf{r}+\mathbf{e}^{\prime\prime}+\mu\lceil\frac{q}{2}\rceil)$. With this smudging $\mathbf{e}^{\prime}\mathbf{r}$ can now be drowned by $\mathbf{e}^{\prime\prime}$ effectively now behaving as a fresh LWE sample in $\mathbf{x}_{0}$ with slightly larger noise. Accordingly, to make the final security proof go through, we design our search-to-decision reduction to also work for samples statistically-close to LWE versus samples close to random, as opposed to the simple case of real LWE versus truly random.